When 200 mL of 0.700 molarity silver nitrate is introduced to 150 mL of 0.500 molarity potassium chromate, how many grams of silver chromate will precipitate? (12 points)




____ AgNO3(aq) + ____K2CrO4(aq) → ____Ag2CrO4(s)+ ____KNO3(aq)





1
Expert's answer
2022-03-28T17:19:04-0400

2 AgNO3(aq) + K2CrO4(aq) → Ag2CrO4(s) + 2KNO3(aq)

CM = n / V

n = CM x V

n (AgNO3) = CM x V = 0.700 x 0.2 = 0.14 mol

n (K2CrO4) = CM x V = 0.500 x 0.15 = 0.075 mol

According to the equation, 2 n (AgNO3(aq)) = n (K2CrO4) = n (Ag2CrO4). Therefore, K2CrO4 is the limiting reactant.

n (Ag2CrO4) = n (K2CrO4) = 0.075 mol

n = m / M

M (Ag2CrO4) = 331.7 g/mol

m (Ag2CrO4) = n x M = 0.075 x 331.7 = 24.9 g